### v2

# Mountain Pass Solutions for an entire semipositone problem involving the Grushin Subelliptic Operator

Abstract
Abstract. For $N\geq 3$ we study the following semipositone problem
$-\Delta_{\gamma}u=g(z)f_{a}(u)\quad\hbox{in $\mathbb{R}^{N}$},$
where $\Delta_{\gamma}$ is the Grushin operator
$\Delta_{\gamma}u(z)=\Delta_{x}u(z)+\lvert x\rvert^{2\gamma}\Delta_{y}u(z)\quad(\gamma\geq 0),$
$g\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ is a positive function, $a>0$ is a parameter and $f_{a}$ is a continuous function on $\mathbb{R}$ that coincides with $f(t)-a$ for $t\in\mathbb{R}^{+}$, where $f$ is a continuous function with subcritical and Ambrosetti-Rabinowitz type growth and which satisfies $f(0)=0$. Depending on the range of $a$, we obtain the existence of positive mountain pass solutions in $D_{\gamma}(\mathbb{R}^{N})$, extending the results found in [1, 7, 8].
1 Introduction
The term semipositone is usually associated to a semilinear elliptic equation of the form
$-\Delta u=f(u)\quad\hbox{in $\Omega$}$
in which the nonlinearity $f$ is such that $f(0)<0$, see [14]. In particular, $u=0$ is not a subsolution of the equation, and it was remarked in [28] that this fact may introduce several difficulties in the analysis of the problem. When the domain $\Omega$ is unbounded the condition $f(0)<0$ is also a source of concern from the viewpoint of variational methods, since non-zero constants are not integrable on $\Omega$.
In literature, second order semipositone problems have been widely studied. In the seminal paper [11], Brown and Shivaji considered the perturbed bifurcation problem $-\Delta u=\lambda(u-u^{3})-\varepsilon$ on a bounded set $\Omega\subset\mathbb{R}^{N}$ with Dirichlet boundary conditions and $\varepsilon,\lambda>0$.
Some years later, Castro and Shivaji (see [13]) studied non-negative solutions for the boundary value problem
$\begin{cases}-u^{\prime\prime}(x)=\lambda f(u(x)),\leavevmode\nobreak\ x\in(0,1)\\ u(0)=0=u(1),\end{cases}$
where $\lambda>0$ is a constant and $f\in C^{2}(\mathbb{R})$ satisfies $f(0)<0$.
Subsequently, in [12] the authors considered positive solutions for a class of multiparameter semipositone problem
$\begin{cases}-\Delta u=\lambda g(u)+\mu f(u)\leavevmode\nobreak\ &\hbox{in $\Omega$},\\ u=0\leavevmode\nobreak\ &\hbox{on $\partial\Omega$},\end{cases}$
with $\lambda>0$, $\mu>0$ and $\Omega$ is a smooth bounded domain in $\mathbb{R}^{N}$, $N\geq 2$. They assume $g(0)>0$ and superlinear while $f(0)<0$ sublinear, and eventually strictly positive. For fixed $\mu$, they establish existence and multiplicity for $\lambda$ small (when the problem becomes semipositone), and nonexistence for $\lambda$ large.
To the best of our knowledge Alves, Holanda, and Santos in [1] were the first to study semipositone problems in the entire space $\mathbb{R}^{N}$, $N\geq 3$, for the classical Laplace operator, described by the equation
$\begin{cases}-\Delta u=h(x)\left(f(u)-a\right)\leavevmode\nobreak\ &\hbox{in $\mathbb{R}^{N}$, }\\ u>0\leavevmode\nobreak\ &\hbox{in $\mathbb{R}^{N}$},\end{cases}$
with $f(0)=0$ and $a>0$. The function $f\in C(\mathbb{R}^{+})$ is a locally Lipschitz function with subcritical growth which satisfies the Ambrosetti-Rabinowitz condition, and the weight $h$ is a continuous positive function bounded by a radial function $P\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$.
Subsequently, many authors extended their result to other operators. In [9], Biswas, Das and Sarkar examined the same problem for a fourth-order operator. Additionally, in [7], Biswas investigated this problem for the fractional Laplacian. In the very recent preprint [8], a semipositone problem with the fractional $p$-Laplace operator was investigated.
In this paper we will be concerned with the following semipositone problem associated with the Grushin operator
$-\Delta_{\gamma}u=g(z)f_{a}(u)\leavevmode\nobreak\ \hbox{in $\mathbb{R}^{N}$},$ (SP)
where the function $g\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ is positive, $a>0$ is a parameter and the function $f_{a}\colon\mathbb{R}\to\mathbb{R}$ is defined as:
$f_{a}(t)=\begin{cases}f(t)-a&\hbox{if $t\geq 0$},\\ -a(t+1)&\hbox{if $t\in(-1,0)$},\\ 0&\hbox{if $t\leq-1$},\end{cases}$
where $f$ is a contiuous function defined on $[0,+\infty)$ is such that $f(0)=0$ and
  1. (f1)
    $\lim_{t\to 0^{+}}\frac{f(t)}{t}=0$, $\lim_{t\to\infty}\frac{f(t)}{t^{\alpha-1}}<+\infty$ for some $\alpha\in(2,2^{*}_{\gamma})$,
  2. (f2)
    (Ambrosetti-Rabinowitz) there exist $\theta>2$ and $t_{0}>0$ such that
    $0<\theta F(t)\leq tf(t)$
    for all $t>t_{0}$, where $F(t)=\int_{0}^{t}f(s)\,\mathrm{d}s$.
For the reader’s convenience, we briefly recall the basic construction of the Baouendi-Grushin operator. We split the euclidean space $\mathbb{R}^{N}$ as $\mathbb{R}^{m}\times\mathbb{R}^{\ell}$, where $m\geq 1$, $\ell\geq 1$ and $m+\ell=N\geq 3$. A generic point $z\in\mathbb{R}^{N}$ can therefore be written as
$z=\left(x,y\right)=\left(x_{1},\ldots,x_{m},y_{1},\ldots,y_{\ell}\right),$
where $x\in\mathbb{R}^{m}$ and $y\in\mathbb{R}^{\ell}$. For a differentiable function $u$ we define the Grushin gradient
$\nabla_{\gamma}u(z)=(\nabla_{x}u(z),\lvert x\rvert^{\gamma}\nabla_{y}u(z))=(\partial_{x_{1}}u(z),\dots,\partial_{x_{m}}u(z),\lvert x\rvert^{\gamma}\partial_{y_{1}}u(z),\dots,\lvert x\rvert^{\gamma}\partial_{y_{\ell}}u(z)).$
The Grushin operator $\Delta_{\gamma}$ is defined by [^1] Some authors use the slightly different definition $\Delta_{\gamma}=\Delta_{x}+\left(1+\gamma\right)^{2}|x|^{2\gamma}\Delta_{y}$.
$\Delta_{\gamma}u(z)=\Delta_{x}u(z)+\lvert x\rvert^{2\gamma}\Delta_{y}u(z),$
where $\Delta_{x}$ and $\Delta_{y}$ are the Laplace operators in the variables $x$ and $y$, respectively. A crucial property of the Grushin operator is that it is not uniformly elliptic in the space $\mathbb{R}^{N}$, since it is degenerate on the subspace $\{0\}\times\mathbb{R}^{\ell}$.
For a generic $\gamma\geq 0$, it is possible to embed $\Delta_{\gamma}$ into a family of operators of the form
$\Delta_{\lambda}=\sum_{i=1}^{N}\partial_{x_{i}}\left(\lambda_{i}^{2}\partial_{x_{i}}\right),$
which have been studied from a geometrical viewpoint in [19, 20, 21]. Here $\lambda_{1},\ldots,\lambda_{N}$ are given functions which satisfy suitable conditions.
Moreover, the Grushin operator falls into the class of $X$-elliptic operators studied in [22, 27]. In fact, the Grushin operator is uniformly $X$-elliptic with respect to the family of vector fields $X=(X_{1},\dots,X_{N})$ defined as
$X_{i}$ $=\frac{\partial}{\partial x_{i}}\quad\hbox{for $i=1,\ldots,m$}$
$X_{m+j}$ $=|x|^{\gamma}\frac{\partial}{\partial y_{j}}\quad\hbox{for $j=1,\ldots,\ell$}.$
With the same notation we can write the Grushin operator $\Delta_{\gamma}$ as sum of the vector fields
$\Delta_{\gamma}=\sum_{j=0}^{N}X_{j}^{2}.$
Note that, when $\gamma$ is an even integer, the $C^{\infty}$-field $X=\left(X_{1},\ldots,X_{N}\right)$ satisfies the Hörmander’s finite rank condition, see [24].
We will look for solutions to (SP) in the function space
$D_{\gamma}(\mathbb{R}^{N})=\left\{u\in L^{2^{*}_{\gamma}}(\mathbb{R}^{N}):\lVert u\rVert_{\gamma}<\infty\right\},$ (1)
where
$\lVert u\rVert_{\gamma}=\left(\int_{\mathbb{R}^{N}}\lvert\nabla_{\gamma}u\rvert^{2}\,{\mathrm{d}z}\right)^{1/2}.$
By definition, a function $u\in D_{\gamma}(\mathbb{R}^{N})$ is a weak solution to (SP) if it satisfies the following variational identity:
$\forall\varphi\in D_{\gamma}(\mathbb{R}^{N}):\quad\int_{\mathbb{R}^{N}}\nabla_{\gamma}u\nabla_{\gamma}\varphi\,{\mathrm{d}z}=\int_{\mathbb{R}^{N}}g(z)f_{a}(u)\varphi\,{\mathrm{d}z}.$
Two functionals $K_{a}$ and $I_{a}$ are defined on $D_{\gamma}(\mathbb{R}^{N})$ by
$K_{a}(u)$ $=\int_{\mathbb{R}^{N}}g(z)F_{a}(u(z))\,{\mathrm{d}z}$
$I_{a}(u)$ $=\frac{1}{2}\lVert u\rVert_{\gamma}^{2}-K_{a}(u).$
It is standard to prove that $K_{a}$ and $I_{a}$ are of class $C^{1}$, and the corresponding Fréchet derivatives are given by
$DK_{a}(u)\colon v\to\int_{\mathbb{R}^{N}}g(z)f_{a}(u)v\,{\mathrm{d}z}\leavevmode\nobreak$
$DI_{a}(u)\colon v\to\int_{\mathbb{R}^{N}}\nabla_{\gamma}u\nabla_{\gamma}v\,{\mathrm{d}z}.$
As a consequence, weak solutions of (SP) are precisely the critical points of $I_{a}$ on $D_{\gamma}(\mathbb{R}^{N})$.
The magnitude of the constant $a$ and the interplay with the function $g$ are crucial. Specifically, we can prove the existence of positive solutions to (SP) when the parameter $a$ is sufficiently small.
We will construct a solution by a mountain pass argument. The positivity of the solution is obtained using the fundamental solution for the Grushin operator.
Our main result can be stated as follows. The distance $d(z)$ appearing in the statement is defined in (31) below.
Theorem 1.1.
Let $f$ satisfy (f1)-(f2) and let $g\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ be positive. Then there exists $a_{1}>0$ such that for each $a\in(0,a_{1})$ (SP) admits a weak solution $u_{a}\in C(\mathbb{R}^{N})\cap D_{\gamma}(\mathbb{R}^{N})$ such that $u_{a}\geq 0$ a.e. in $\{0\}\times\mathbb{R}^{\ell}$. If, in addition,
  1. (f3)
    $f$ is locally Lipschitz,
  1. (g1)
    for all $z\in(\{0\}\times\mathbb{R}^{\ell})\setminus\{(0,0)\}$,
    $d(z)^{N_{\gamma}-2}\int_{\mathbb{R}^{N}}\frac{g(\xi)}{d(z-\xi)^{N_{\gamma}-2}}\mathrm{d}\xi<\infty,$
then there exists $\tilde{a}\in(0,a_{1})$ such that $u_{a}>0$ a.e. in $\{0\}\times\mathbb{R}^{\ell}$ for every $a\in(0,\tilde{a})$.
Remark 1.2.
Actually, standard regularity results (see for instance [31] and the references there-in for the case of the Heisenberg group, but the proof extends to our setting) ensure that $u\in C^{1,\beta}(\mathbb{R}^{N})$ with $10<\beta<$ 1.
To the best of our knowledge, there are no results in the literature about semipositone problems for degenerate equations of Baouendi-Grushin type. Before concluding this introduction, we would like to comment about the positivity in the subspace $\{0\}\times\mathbb{R}^{\ell}$. In some sense this is an expected result, since the restriction of $\Delta_{\gamma}$ to this subspace is a uniformly elliptic operator for which all the standard tools are available. Globally, the Grushin operator lacks a strong maximum principle for weak solutions and $W^{2,p}$-estimates which are used in order to derive additional regularity for weak solutions. We refer to Remark 6.8 for more details.
2 Some inequalities for $f$
Remark 2.1.
In the rest of the paper we will denote by $C(\ldots)$ a positive constant which depends on the variables between brackets.
Let us now identify the upper and lower bounds of $f$, $f_{a}$ and their primitives. Clearly, the function $f_{a}$ is continuous and the function $F_{a}$ defined by
$F_{a}(t)=\begin{cases}F(t)-at\leavevmode\nobreak\ &\hbox{for $t\geq 0$},\\ -\frac{at^{2}}{2}-at\leavevmode\nobreak\ &\hbox{for $t\in(-1,0)$},\\ \frac{a}{2}\leavevmode\nobreak\ &\hbox{for $t\leq-1$}\end{cases}$
satisfies $F_{a}^{\prime}=f_{a}$ and $F_{a}(0)=0$. The assumption
$\lim_{t\to 0^{+}}\frac{f(t)}{t}=0$
implies that for every $\varepsilon>0$, there exists $t_{1}(\varepsilon)>0$ such that $f(t)<\varepsilon t$ for $t\in(0,t_{1}(\varepsilon))$. Similarly, assumption
$\lim_{t\to\infty}\frac{f(t)}{t^{\alpha-1}}<+\infty$
implies
$f(t)\leq C(f,t_{1}(\varepsilon))t^{\alpha-1}\quad\hbox{for $t\geq t_{1}(\varepsilon)$}$
for some positive constant $C(f,t_{1}(\varepsilon))$. Hence, we have the following bounds for every $\alpha\in(2,2^{*}_{\gamma}]$:
$\lvert f_{a}(t)\rvert\leq\varepsilon\lvert t\rvert+C(f,t_{1}(\varepsilon))\lvert t\rvert^{\alpha-1}+a,\quad F_{a}(t)\leq\varepsilon\lvert t\rvert^{2}+C(f,t_{1}(\varepsilon))\lvert t\rvert^{\alpha}+a\lvert t\rvert\quad\hbox{for $t\in\mathbb{R}$}.$ (2)
By (f2), the function $t\mapsto\frac{F(t)}{t^{\theta}}$ is non decreasing for $t>t_{0}$, so there exist two positive constants $M_{1}$ and $M_{2}$ such that
$F(t)\geq M_{1}t^{\theta}-M_{2}\quad\hbox{for all $t\in\mathbb{R}$}.$ (3)
Remark 2.2 (Ambrosetti-Rabinowitz condition on $f_{a}$).
For $t>t_{0}$, assumption (f2) yields
$\theta F_{a}(t)=\theta F(t)-\theta at\leq f(t)t-at=tf_{a}(t).$
For $t\in[0,t_{0}]$, by continuity of $F$, there exists a constant $M>0$ independent of $a$, such that
$\theta F_{a}(t)=\theta F(t)-\theta at\leq M-at\leq M-at+f(t)t\leq M+tf_{a}(t)$
For $t\in[-1,0]$,
$\theta F_{a}(t)-tf_{a}(t)=\left(-\frac{a\theta}{2}+a\right)t^{2}-\left(\theta-1\right)at\leq ca$
where $c=\max_{t\in[-1,0]}\left\{\left(-\frac{\theta}{2}+1\right)t^{2}-\left(\theta-1\right)t\right\}$ does not depend on $a$.
For $t<-1$,
$\theta F_{a}(t)=\frac{\theta a}{2}\leq\frac{\theta a}{2}+f_{a}(t)t$
Choosing $C=\max\left\{c,\frac{\theta}{2}\right\}$ we conclude that the following Ambrosetti-Rabinowitz condition with correction terms holds for $f_{a}$:
$\theta F_{a}(t)\leq tf_{a}(t)+Ca+M,\quad\hbox{for all $t\in\mathbb{R}$ and $a>0$}.$ (4)
3 A functional setting for the Grushin operator
The space $D_{\gamma}(\mathbb{R}^{N})$ defined in (1) can be also characterized as the completion of $C_{0}^{\infty}(\mathbb{R}^{N})$ with respect to the norm
$\lVert u\rVert_{\gamma}=\left(\int_{\mathbb{R}^{N}}\lvert\nabla_{\gamma}u\rvert^{2}\,{\mathrm{d}z}\right)^{1/2},$
where
$2_{\gamma}^{*}=\frac{2N_{\gamma}}{N_{\gamma}-2}$
is the Sobolev critical exponent associated to $N_{\gamma}=m+(1+\gamma)\ell$, the homogeneous dimension associated to the decomposition $N=m+\ell$. Since the Sobolev-Gagliardo-Niremberg inequality proved in [29, Lemma 5.1]
$\lVert u\rVert_{2^{*}_{\gamma}}\leq C\lVert u\rVert_{\gamma},$
holds for all $u\in D_{\gamma}(\mathbb{R}^{N})$ for some positive constant $C=C(m,k,\gamma)$, we have that the embedding
$D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{2^{*}_{\gamma}}(\mathbb{R}^{N})$ (5)
is continuous. The homogeneous space $D_{\gamma}(\mathbb{R}^{N})$ is also a Hilbert space with the inner product [^2] For the sake of simplicity, we omit any scalar multiplication sign between the two vectors $\nabla_{\gamma}u$ and $\nabla_{\gamma}v$.
$\braket{u,v}_{\gamma}=\int_{\mathbb{R}^{N}}\nabla_{\gamma}u\nabla_{\gamma}v\,\,{\mathrm{d}z}.$
Given $q\in[1,\infty)$, with $L^{q}(\mathbb{R}^{N},\lvert g\rvert)$ we denote the class of real-valued Lebesgue measurable functions $u$ such that
$\int_{\mathbb{R}^{N}}\lvert g(z)\rvert\lvert u(z)\rvert^{q}\,{\mathrm{d}z}<+\infty.$
By the assumptions on $g$ and (5), we have that the embedding $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{q}(\mathbb{R}^{N},\lvert g\rvert)$ is continuous for $q\in[1,2^{*}_{\gamma}]$.
The following lemma states a compact embedding of the solution space for (SP) into the weighted Lebesgue space $L^{q}(\mathbb{R},\lvert g\rvert)$.
Lemma 3.1.
Let $q\in[1,2^{*}_{\gamma})$. The embedding $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{q}(\mathbb{R}^{N},\lvert g\rvert)$ is compact.
Proof.
Let $\{u_{j}\}$ be a bounded sequence in $D_{\gamma}(\mathbb{R}^{N})$. We can suppose that $u_{j}\rightharpoonup 0$. For each $R>0$, we have the continuous embedding $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow\mathscr{H}_{\gamma}(B_{R}(0))$ [^3] For an open bounded set $\Omega\subset\mathbb{R}^{N}$, we denote by $\mathscr{H}_{\gamma}(\Omega)$ the completion of $C^{1}_{0}(\Omega)$ with respect to the norm $\lVert u\rVert_{\gamma}=\left(\int_{\Omega}\lvert\nabla_{\gamma}u\rvert^{2}\,{\mathrm{d}z}\right)^{1/2}.$ Since the embedding $\mathscr{H}_{\gamma}(B_{R}(0))\hookrightarrow L^{q}(B_{R}(0))$ is compact, see [26, Proposition 3.2], it follows that $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{q}(B_{R}(0))$ is a compact embedding as well. Hence,
$u_{j}(x)\to 0\leavevmode\nobreak\ \hbox{a.e. in $\mathbb{R}^{N}$}$
for some subsequence. As $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{2^{*}_{\gamma}}(\mathbb{R}^{N})$ is a continuous embedding, we have that $\{\lvert u_{j}\rvert^{q}\}$ is a bounded sequence in $L^{2^{*}_{\gamma}/q}(\mathbb{R}^{N})$. By the Brezis-Lieb Lemma, [10, Theorem 1], up to a subsequence if necessary
$\lvert u_{j}\rvert^{q}\rightharpoonup 0\leavevmode\nobreak\ \hbox{in $L^{2^{*}_{\gamma}/q}(\mathbb{R}^{N})$},$
or equivalently,
$\int_{\mathbb{R}^{N}}\lvert u_{j}\rvert^{q}\varphi\,{\mathrm{d}z}\to 0,\leavevmode\nobreak\ \forall\varphi\in L^{p}(\mathbb{R}^{N}),$
where $p$ is the Hölder conjugate of $2^{*}_{\gamma}/q$. From the assumptions on $g$ we know that $g\in L^{r}(\mathbb{R}^{N})$ for all $r\geq 1$. It follows that
$\int_{\mathbb{R}^{N}}g(z)\lvert u_{j}\rvert^{q}\,{\mathrm{d}z}\to 0.$
This shows that $u_{j}\to 0$ in $L^{q}(\mathbb{R}^{N},\lvert g\rvert)$, concluding the proof. ∎
4 Existence of solutions for (SP)
In the next lemma, depending on the values of $a$, we will establish the mountain pass geometry for the functional $I_{a}$.
Lemma 4.1.
Let $N\geq 3$. Let $f$ satisfies (f1), (f2) and $g\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ be positive. Then the following hold:
  1. (i)
    There exists $r>0$ such that if $\rho\in(0,r)$ and $\lVert u\rVert_{\gamma}=\rho$, then there are $\delta=\delta(\rho)>0$ and $a_{1}=a_{1}(\rho)>0$ such that if $I_{a}(u)\geq\delta$ for all $a\in(0,a_{1})$.
  2. (ii)
    There exists $v\in D_{\gamma}(\mathbb{R}^{N})$ such that $\lVert v\rVert_{\gamma}>\rho$ and $I_{a}(v)<0$, for all $a\in(0,a_{1})$.
Proof.
  1. (i)
    Using (2), we estimate
    $\int_{\mathbb{R}^{N}}g(z)F_{a}(u)\,{\mathrm{d}z}$ $\leq\int_{\mathbb{R}^{N}}g(z)\left(\varepsilon\lvert u\rvert^{2}+C(f,\varepsilon)\lvert u\rvert^{\alpha}+a\lvert u\rvert\right)\,{\mathrm{d}z}$
    $=\varepsilon\int_{\mathbb{R}^{N}}g(z)\lvert u\rvert^{2}\,{\mathrm{d}z}+C(f,\varepsilon)\int_{\mathbb{R}^{N}}g(z)\lvert u\rvert^{\alpha}\,{\mathrm{d}z}+a\int_{\mathbb{R}^{N}}g(z)\lvert u\rvert\,{\mathrm{d}z}$
    $\leq\varepsilon C_{1}\lVert u\rVert_{\gamma}^{2}+C_{2}\lVert u\rVert_{\gamma}^{\alpha}+aC_{3}\lVert u\rVert_{\gamma},$
    where $C_{1}$, $C_{2}$ and $C_{3}$ are the embedding constants of $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{q}(\mathbb{R}^{N},\lvert g\rvert)$, see Lemma 3.1. In particular, for $\lVert u\rVert_{\gamma}=\rho$,
    $I_{a}(\rho)$ $\geq\frac{1}{2}\rho^{2}-\varepsilon C_{1}\rho^{2}-C_{2}\rho^{\alpha}-aC_{3}\rho$
    $=\rho^{2}\left(\frac{1}{2}-\varepsilon C_{1}-C_{2}\rho^{\alpha-2}\right)-aC_{3}\rho.$ (6)
    We choose $\varepsilon<(2C_{1})^{-1}$. Then we write $I_{a}(u)\geq A(\rho)-aC_{3}\rho$, where $A(\rho)=C\rho^{2}\left(1-\tilde{C}\rho^{\alpha-2}\right)$, with $C$, $\tilde{C}$ independent of $a$. Let $r$ be the first non-trivial zero of $A$. For $\rho<r$, we fix $a_{1}\in\left(0,\frac{A(\rho)}{C_{3}\rho}\right)$ and $\delta=A(\rho)-a_{1}C_{3}\rho$. Thus, by (6) we get $I_{a}(u)\geq\delta$ for every $a\in(0,a_{1})$.
  2. (ii)
    Fix a function
    $\varphi\in C^{\infty}_{0}(\mathbb{R}^{N})\setminus\{0\},\quad\hbox{with $\varphi\geq 0$\leavevmode\nobreak\ and $\lVert\varphi\rVert_{\gamma}=1$.}$
    Notice that for all $t>0$,
    $I_{a}(t\varphi)=\frac{1}{2}\int_{\Omega}\lvert\nabla_{\gamma}t\varphi\rvert^{2}\,{\mathrm{d}z}-\int_{\Omega}g(z)F_{a}(t\varphi)\,{\mathrm{d}z}$
    $=\frac{t^{2}}{2}-\int_{\Omega}g(z)F(t\varphi)\,{\mathrm{d}z}+a\int_{\Omega}g(z)t\varphi\,{\mathrm{d}z}$
    where $\Omega=\operatorname{supp}\varphi$. By (3),
    $I_{a}(t\varphi)\leq\frac{t^{2}}{2}-M_{1}t^{\theta}\int_{\Omega}g(z)\lvert\varphi\rvert^{\theta}\,{\mathrm{d}z}+ta\lVert g\rVert_{\infty}+M_{2}\lVert g\rVert_{1}.$
    Since $\theta>2$ and $a\in(0,a_{1})$, we can fix $t_{0}>1$ large enough so that $I_{a}(v)<0$, where $v=t_{0}\varphi\in D_{\gamma}(\mathbb{R}^{N})$.
As a further step we prove the Palais-Smale condition of $I_{a}$. Recall that from the estimate found in (4), for a fixed $a>0$,
$\theta F_{a}(t)\leq f_{a}(t)t+M\quad\hbox{for all $t\in\mathbb{R}$},$ (7)
for a suitable positive $M$.
Lemma 4.2.
The functional $I_{a}$ satisfies the Palais-Smale condition for all $a>0$.
Proof.
Let $\{u_{j}\}_{j}$ be a Palais-Smale sequence in $D_{\gamma}(\mathbb{R}^{N})$. First of all we claim that
the sequence $\{u_{j}\}$ is bounded in $D_{\gamma}(\mathbb{R}^{N})$. (8)
Since $\{u_{j}\}_{j}$ is a Palais-Smale sequence, there exists $j_{0}\in\mathbb{N}$ such that
$\left|\left\langle I^{\prime}_{a}(u_{j}),u_{j}\right\rangle\right|\leq\lVert u_{j}\rVert_{\gamma}$
for $j\geq j_{0}$, and there exists $k>0$ such that for any $j\in\mathbb{N}$
$\lvert I_{a}(u_{j})\rvert\leq k.$
So
$-\lVert u_{j}\rVert_{\gamma}-\lVert u_{j}\rVert_{\gamma}^{2}\leq-\int_{\mathbb{R}^{N}}g(z)f_{a}(u_{j})u_{j}\,{\mathrm{d}z}$ (9)
and
$\frac{1}{2}\lVert u_{j}\rVert_{\gamma}^{2}-\int_{\mathbb{R}^{N}}g(z)F_{a}(u_{j})\,{\mathrm{d}z}\leq k$ (10)
for all $j\geq j_{0}$. From (7) and (10),
$\frac{1}{2}\lVert u_{j}\rVert_{\gamma}^{2}-\frac{1}{\theta}\int_{\mathbb{R}^{N}}g(z)f_{a}(u_{j})u_{j}\,{\mathrm{d}z}-\frac{1}{\theta}M\lVert g\rVert_{1}\leq k$ (11)
for all $j\geq j_{0}$. Thereby, by (9) and (11),
$\frac{1}{2}\lVert u_{j}\rVert_{\gamma}^{2}-\frac{1}{\theta}\lVert u_{j}\rVert_{\gamma}-\frac{1}{\theta}\lVert u_{j}\rVert_{\gamma}^{2}\leq k+\frac{1}{\theta}M\lVert g\rVert_{1},$
or equivalently,
$\left(\frac{1}{2}-\frac{1}{\theta}\right)\lVert u_{j}\rVert_{\gamma}^{2}-\frac{1}{\theta}\lVert u_{j}\rVert_{\gamma}\leq k+\frac{1}{\theta}M\lVert g\rVert_{1},$
for $j$ large enough, so (8) is proved.
Up to a subsequence, still denoted by $\{u_{j}\}_{j}$, there exists $u\in D_{\gamma}(\mathbb{R}^{N})$ such that
$u_{j}\rightharpoonup u\leavevmode\nobreak\ \hbox{ in $D_{\gamma}(\mathbb{R}^{N})$}$
and
$u_{j}\to u\quad\hbox{a.e. in\leavevmode\nobreak\ $\mathbb{R}^{N}$}.$
From (2) we can estimate
$\lvert g(z)f_{a}(u_{j})\left(u_{j}-u\right)\rvert\leq Cg(z)\lvert u_{j}-u\rvert\left(\lvert u_{j}\rvert+\lvert u_{j}\rvert^{\alpha-1}+a\right)$
for some constant $C>0$ independent of $a$. Now we we claim that
$\int_{\mathbb{R}^{N}}\left(g(z)\lvert u_{j}-u\rvert\left(\lvert u_{j}\rvert+\lvert u_{j}\rvert^{\alpha-1}+a\right)\,\right){\mathrm{d}z}\to 0\leavevmode\nobreak\ \hbox{as \leavevmode\nobreak\ $j\to+\infty$}.$ (12)
In fact, we will only show the limit
$\int_{\mathbb{R}^{N}}\left(g(z)\lvert u_{j}-u\rvert\lvert u_{j}\rvert^{\alpha-1}\,\right){\mathrm{d}z}\to 0\leavevmode\nobreak\ \hbox{as \leavevmode\nobreak\ $j\to+\infty$}.$ (13)
because the limits involving the other terms follow with the same idea. The sequence $\{\lvert u_{j}\rvert^{\alpha-1}\}_{j}$ is bounded in $L^{\frac{2^{*}_{\gamma}}{\alpha-1}}(\mathbb{R}^{N})$, as (8) holds. Moreover, applying Lemma 3.1 with $q=\frac{2^{*}_{\gamma}}{2^{*}_{\gamma}-(\alpha-1)}$, we get
$u_{j}\to u\in L^{\frac{2^{*}_{\gamma}}{2^{*}_{\gamma}-(\alpha-1)}}(\mathbb{R}^{N},\lvert g\rvert).$ (14)
Then, by Hölder’s inequality with conjugate pair $\left(\frac{2^{*}_{\gamma}}{\alpha-1},\frac{2^{*}_{\gamma}}{2^{*}_{\gamma}-\left(\alpha-1\right)}\right)$,
$\int_{\mathbb{R}^{N}}g(z)\lvert u_{j}-u\rvert\lvert u_{j}\rvert^{\alpha-1}\,{\mathrm{d}z}\leq\lVert g\left(u_{j}-u\right)\rVert_{{2^{*}_{\gamma}}/{2^{*}_{\gamma}-\left(\alpha-1\right)}}\lVert u_{j}\rVert_{2^{*}_{\gamma}}^{\alpha-1}.$
Now, by (14) we get (13).
A consequence of (12) is the limit
$\int_{\mathbb{R}^{N}}g(z)f_{a}(u_{j})(u_{j}-u)\,{\mathrm{d}z}\to 0$
and so
$\int_{\mathbb{R}^{N}}\nabla_{\gamma}u_{j}\nabla_{\gamma}\left(u_{j}-u\right)\,{\mathrm{d}z}\to 0.$ (15)
The weak convergence $u_{j}\to u$ in $D_{\gamma}(\mathbb{R}^{N})$ yields
$\int_{\mathbb{R}^{N}}\nabla_{\gamma}u\nabla_{\gamma}\left(u_{j}-u\right)\,{\mathrm{d}z}\to 0.$ (16)
From (15) and (16),
$\int_{\mathbb{R}^{N}}\lvert\nabla_{\gamma}u_{j}-\nabla_{\gamma}u\rvert^{2}\,{\mathrm{d}z}\to 0,$
concluding the proof. ∎
Now we are in a position to apply the Mountain Pass Theorem [3, Theorem 2.1], which ensures the existence and uniform boundedness of solutions to the (SP) problem.
Theorem 4.3.
Let $a_{1}>0$ be given in Lemma 4.1. Then for each $a\in(0,a_{1})$, (SP) has a solution $u_{a}\in D_{\gamma}(\mathbb{R}^{N})$. Moreover, there exists $C>0$ such that $\lVert u\rVert_{\gamma}\leq C$ for all $a\in(0,a_{1})$.
Proof.
Consider $a_{1}$, $\delta$ and $v$ as given in Lemma 4.1. For $a\in(0,a_{1})$, using Lemma 4.1 and 4.2, the hypotheses of the mountain pass theorem [3, Theorem 2.1] are verified. So there exists a non-trivial critical point $u_{a}\in D_{\gamma}(\mathbb{R}^{N})$ of $I_{a}$ satisfying
$I_{a}(u_{a})=c_{a}=\inf_{\gamma\in\Gamma_{v}}\max_{t\in[0,1]}I_{a}(\gamma(t))\geq\delta\leavevmode\nobreak\ \hbox{and $I^{\prime}_{a}(u_{a})=0$},$ (17)
where $\Gamma_{v}=\{\gamma\in C\left([0,1],D_{\gamma}(\mathbb{R}^{N})\right)\mid\gamma(0)=0,\gamma(1)=v\}$. Thus, $u_{a}$ is a non-trivial solution of (SP). To prove the uniform boundedness of $u_{a}$ in $D_{\gamma}(\mathbb{R}^{N})$, we first show that the set $\{I_{a}(u_{a})\mid a\in(0,a_{1})\}$ is uniformly bounded. Considering $t_{1}$ and $\varphi$ as in Lemma 4.1-(ii), Let us introduce a path $\tilde{\gamma}\colon[0,1]\to D_{\gamma}(\mathbb{R}^{N})$ defined as $\tilde{\gamma}(\sigma)=\sigma v$, where $v=t\varphi$ for some $t>t_{1}$. From (17) we have that
$I_{a}(u_{a})\leq\max_{\sigma\in[0,1]}I_{a}(\tilde{\gamma}(\sigma))=\max_{\sigma\in[0,1]}I_{a}(\sigma t\varphi).$ (18)
Now
$I_{a}(\sigma t\varphi)$ $\leq\frac{\sigma^{2}t^{2}}{2}\lVert\varphi\rVert_{\gamma}^{2}-M_{1}\sigma^{\theta}t^{\theta}\int_{\mathbb{R}^{N}}g(z)\left(\varphi(z)\right)^{\theta}\,{\mathrm{d}z}+M_{2}\int_{\mathbb{R}^{N}}g(z)\,{\mathrm{d}z}+a\sigma t\int_{\mathbb{R}^{N}}g(z)\varphi(z)\,{\mathrm{d}z}$
$\leq\frac{t^{2}}{2}+M_{2}\lVert g\rVert_{1}+a_{1}tC_{1}\lVert\varphi\rVert_{\gamma}.$ (19)
From (18) and (19), there exists $C=C(N_{\gamma},M_{2},g,a_{1})$ such that
$I_{a}(u_{a})\leq C,\leavevmode\nobreak\ \hbox{for all $a\in(0,a_{1})$.}$ (20)
Now we are going to prove that the solutions $u_{a}$ are uniformly bounded in $D_{\gamma}(\mathbb{R}^{N})$ with respect to $a\in(0,a_{1})$. From (18) we know that
$\lVert u_{a}\rVert_{\gamma}^{2}-\int_{\mathbb{R}^{N}}g(z)f_{a}(z)u_{a}\,{\mathrm{d}z}=0.$ (21)
Also, by (20) we have
$\frac{1}{2}\lVert u_{a}\rVert_{\gamma}^{2}-\int_{\mathbb{R}^{N}}g(z)F_{a}(u_{a})\,{\mathrm{d}z}\leq C.$ (22)
Now first multiplying (21) by $\frac{1}{\theta}$ and then subtracting into (22) gives the following
$\left(\frac{1}{2}-\frac{1}{\theta}\right)\lVert u_{a}\rVert_{\gamma}^{2}+\int_{\mathbb{R}^{N}}g(z)\left(\frac{1}{\theta}f_{a}(u_{a})u_{a}-F_{a}(u_{a})\right)\,{\mathrm{d}z}\leq C,$
and combining the above with (7) yields
$\left(\frac{1}{2}-\frac{1}{\theta}\right)\lVert u_{a}\rVert_{\gamma}^{2}-\frac{1}{\theta}M_{3}\lVert g\rVert_{1}\leq C.$
So, there exists $C$ such that $\lVert u_{a}\rVert_{\gamma}\leq C$ for every $a\in(0,a_{1})$. ∎
5 Properties of the solutions
Our next result establishes that the solution $u_{a}$ belongs to $L^{\infty}(\mathbb{R}^{N})$ if the parameter $a$ is sufficiently small. Using the iterative argument of [32, Lemma B.3] we prove that the solution $u_{a}\in L^{q}(\mathbb{R}^{N})$, for every $q\in[2^{*}_{\gamma},\infty)$, provided $a$ is small enough. Moreover the argument in [30, Theorem 3.1] can be applied to establish the boundedness of $u_{a}$.
Lemma 5.1.
There exists $a_{2}\in(0,a_{1})$ such that $u_{a}\in L^{\infty}(\mathbb{R}^{N})\cap C(\mathbb{R}^{N})$ for all $a\in(0,a_{2})$.
Proof.
In order to prove the lemma, it is enough to show that for any sequence $a_{j}\to 0$, the sequence of solutions $u_{j}=u_{a_{j}}$ possesses a subsequence, still denoted by itself, which is bounded in $L^{\infty}(\mathbb{R}^{N})$. By Lemma 4.3, the sequence $\{u_{j}\}$ is bounded in $D_{\gamma}(\mathbb{R}^{N})$, then for some subsequence, there is $u\in D_{\gamma}(\mathbb{R}^{N})$ such that
$u_{j}\rightharpoonup u\quad\hbox{in $D_{\gamma}(\mathbb{R}^{N})$}$
and
$u_{j}\to u\quad\hbox{a.e. in $\mathbb{R}^{N}$}.$
With the same approach explored in the proof of Lemma 4.2, we have that
$u_{j}\to u\quad\hbox{in $D_{\gamma}(\mathbb{R}^{N})$}.$
Consequently,
$\ u_{j}\to u\quad\hbox{in $L^{2^{*}_{\gamma}}(\mathbb{R}^{N})$}$ (23)
and for some subsequence, there is $h\in L^{2^{*}_{\gamma}}(\mathbb{R}^{N})$ such that
$\lvert u_{j}(z)\rvert\leq h(z),\quad\hbox{a.e. in $\mathbb{R}^{N}$ and $\forall j\in\mathbb{N}$}.$
Setting the function
$V_{j}(z)=g(z)\frac{\left(1+\lvert u_{j}\rvert^{\frac{N_{\gamma}+2}{N_{\gamma}-2}}\right)}{1+\lvert u_{j}\rvert},$
it follows that
$\lvert V_{j}\rvert\leq g(1+\lvert u_{j}\rvert^{4/{N_{\gamma}-2}})\leq g(1+\lvert h\rvert^{4/{N_{\gamma}-2}}),\quad\forall j\in\mathbb{N}.$ (24)
By the assumption on the function $g$, $V_{j}\in L^{p}(\mathbb{R}^{N})$, for every $p\in[1,N_{\gamma}/2]$. Thereby, the Dominated Convergence Theorem with (23) and (24) implies that $V_{j}\to V$ in $L^{p}(\mathbb{R}^{N})$, for some $V\in L^{p}(\mathbb{R}^{N})$, $p\in[1,2^{*}_{\gamma}]$. So there exist a constant $C>0$ and $j_{0}\in\mathbb{N}$ such that
$\lvert g(z)f_{a_{j}}(u_{j})\rvert\leq CV(z)\left(1+\lvert u_{j}\rvert\right),\quad\forall j\geq j_{0}.$
In the rest of the proof we will assume $C=1$ to save notation.
Step 1. Given $q\in[2^{*}_{\gamma},\infty)$, there exists a constant $K_{q}$ such that $\lVert u_{j}\rVert_{q}\leq K_{q}$ for all $j\geq j_{0}$.
Let us prove that if $u_{j}\in L^{q}(\mathbb{R}^{N})$ for some $q\geq 2^{*}_{\gamma}$, then $u_{j}\in L^{\frac{2^{*}_{\gamma}}{2}q}$. Let $L\geq 0$ and $G(u_{j})=G_{L}(u_{j})=u_{j}\min\{{u_{j}^{(q/2)-1},L}\}$ and define
$F(u_{j})=F_{L}(u_{j})=\int_{0}^{u_{j}}\lvert G^{\prime}(t)\rvert^{2}\,{\mathrm{d}t}\in D_{\gamma}(\mathbb{R}^{N}).$
First of all, observe that
$\int_{\mathbb{R}^{N}}\nabla_{\gamma}F(u_{j})\nabla_{\gamma}u_{j}\,{\mathrm{d}z}=\int_{\mathbb{R}^{N}}\lvert\nabla_{\gamma}G(u_{j})\rvert^{2}\,{\mathrm{d}z}\leavevmode\nobreak\ \hbox{for every $u_{j}\in D_{\gamma}(\mathbb{R}^{N})$}.$ (25)
and the following growth condition holds
$u_{j}F(u_{j})\leq C_{q}G(u_{j})^{2}\leq C_{q}u_{j}^{q},$
with a constant $C_{q}$ that can be explicitly given by $C_{q}=\frac{q^{2}}{4(q-1)}$.
Using $F(u_{j})$ as a test function in (SP) we get
$\int_{\mathbb{R}^{N}}\nabla_{\gamma}F(u_{j})\nabla_{\gamma}u_{j}\,{\mathrm{d}z}\leq\int_{\mathbb{R}^{N}}V(z)(1+\lvert u_{j}\rvert)u_{j}\min\{u_{j}^{q-2},L^{2}\}\,{\mathrm{d}z}\\ \leq\int_{\mathbb{R}^{N}}V(z)(1+2\lvert u_{j}\rvert^{2})\min\{u_{j}^{q-2},L^{2}\}\,{\mathrm{d}z}\\ \leq\int_{\mathbb{R}^{N}}V(z)\,{\mathrm{d}z}+3\int_{\mathbb{R}^{N}}V(z)\lvert u_{j}\rvert^{2}\min\{u_{j}^{q-2},L^{2}\}\,{\mathrm{d}z}.$ (26)
Moreover, if $q\in L^{q}(\mathbb{R}^{N})$, then for any $K\geq 1$ by (25) and (26) there holds
$\int_{\mathbb{R}^{N}}\lvert\nabla_{\gamma}G(u_{j})\rvert^{2}\,{\mathrm{d}z}\leq c+3\int_{\mathbb{R}^{N}}V(z)\lvert u_{j}\rvert^{2}\min\{u_{j}^{q-2},L^{2}\}\,{\mathrm{d}z}\\ \leq c+3K\int_{\mathbb{R}^{N}}\lvert u_{j}\rvert^{2}\min\{u_{j}^{q-2},L^{2}\}\,{\mathrm{d}z}+3\int_{\{z\in\mathbb{R}^{N}\mid V(z)\geq K\}}\lvert u_{j}\rvert^{2}\min\{u_{j}^{q-2},L^{2}\}\,{\mathrm{d}z}\\ \leq c(1+K)+c\left(\int_{\{z\in\mathbb{R}^{N}\mid V(z)\geq K\}}V(z)^{N_{\gamma}/2}\,{\mathrm{d}z}\right)^{N_{\gamma}/2}\times\left(\int_{\mathbb{R}^{N}}\lvert G(u_{j})\rvert^{\frac{2N_{\gamma}}{N_{\gamma}-2}}\,{\mathrm{d}z}\right)^{\frac{N_{\gamma}-2}{N_{\gamma}}}\\ \leq c(1+K)+c\varepsilon(K)\int_{\mathbb{R}^{N}}\lvert\nabla_{\gamma}G(u_{j})\rvert^{2}\,{\mathrm{d}z},$
with some constants $c$ depending on the $L^{q}(\mathbb{R}^{N})$ norm of $u_{j}$ and
$\varepsilon(K)=\left(\int_{\{z\in\mathbb{R}^{N}\mid V(z)\geq K\}}V(z)^{N_{\gamma}/2}\right)^{2/N_{\gamma}}\to 0\quad\hbox{as $K\to\infty$}.$
Fix $K$ such that $\varepsilon(K)\leq 1/2$, for this choice of $K$ we may conclude that
$\int_{\mathbb{R}^{N}}\lvert\nabla_{\gamma}G(u_{j})\rvert^{2}\,{\mathrm{d}z}$
remains uniformly bounded in $L$. Hence we may let $L\to\infty$ to derive that $u_{j}\in L^{\frac{2^{*}_{\gamma}}{2}q}(\mathbb{R}^{N})$. We can conclude our argument by iterating the process.
Step 2. There exists $C>0$ such that $\lVert u_{j}\rVert_{\infty}\leq C$ for all $j\geq j_{0}$.
Since, by the previous Step, $u_{j}\in L^{q}(\mathbb{R}^{N})$ for all $q\in[1,\infty)$ then also $V\in L^{q}(\mathbb{R}^{N})$ for every $q\in[1,\infty)$. Let $t_{0}>N_{\gamma}/2$, we construct a sequence of $L^{q_{k}}(\mathbb{R}^{N})$ norms of $u_{j}$, with $q_{k}\to\infty$, which are uniformly bounded by the $L^{q_{0}}(\mathbb{R}^{N})$ norm of $u_{j}$, where $q_{0}=t_{0}^{\prime}2^{*}_{\gamma}$ and $t_{0}^{\prime}$ is such that $11/t_{0}+1/t_{0}^{\prime}=$ 1. As before, using again the test function $F(u_{j})$, we estimate the right-hand side of (SP) as follows
$\int_{\mathbb{R}^{N}}V(z)(1+\lvert u_{j}\rvert)F(u_{j})\,{\mathrm{d}z}\leq\lVert V\rVert_{t_{0}}\lVert u_{j}F(u_{j})\rVert_{t_{0}^{\prime}}+C_{q}\lVert V\rVert_{\tilde{t}_{0}}\lVert u_{j}\rVert_{qt_{0}^{\prime}}^{q-1}\\ \leq C_{q}\lVert u_{j}\rVert_{qt_{0}^{\prime}}^{q}\left(\lVert V\rVert_{t_{0}}+\lVert V\rVert_{\tilde{t}_{0}}\lVert u_{j}\rVert_{qt_{0}^{\prime}}^{-1}\right)=\tilde{C}C_{q}\lVert V\rVert_{t_{0}}\lVert u_{j}\rVert_{qt_{0}^{\prime}}^{q}$ (27)
for some $\tilde{t}_{0}>N_{\gamma}/2$ and the constant $\tilde{C}=\tilde{C}(q,V,u_{j})$. Hence, by (25) and (27), we obtain
$\int_{\mathbb{R}^{N}}\lvert\nabla_{\gamma}G(u)\rvert^{2}\,{\mathrm{d}z}\leq\tilde{C}C_{q}\lVert V\rVert_{t_{0}}\lVert u_{j}\rVert_{qt_{0}^{\prime}}^{q}.$
By Sobolev inequality and letting $L\to\infty$, we get
$\lVert u_{j}\rVert_{\frac{2^{*}_{\gamma}}{2}q}^{q}\leq CC_{q}\lVert V\rVert_{t_{0}}\lVert u_{j}\rVert_{qt_{0}^{\prime}}^{q}.$
Let $\delta=\frac{2^{*}_{\gamma}}{2t_{0}^{\prime}}>1$. With this notation, the preceding estimate can be rewritten as
$\lVert u_{j}\rVert_{\delta qt_{0}^{\prime}}\leq\left(CC_{q}\right)^{1/q}\lVert V\rVert_{t_{0}}^{1/q}\lVert u_{j}\rVert_{qt_{0}^{\prime}}.$
Now, define $q_{0}=2^{*}_{\gamma}t_{0}^{\prime}$ and $q_{k}=\delta^{k}q_{0}$ for $k\geq 1$. Observing that $C_{q}\leq Cq$, we obtain
$\lVert u_{j}\rVert_{q^{k}}\leq\left(\prod_{i=0}^{k-1}\left(Cq_{i}\right)^{1/q_{i}}\right)\lVert V\rVert_{t_{0}}^{\sum_{i=0}^{k-1}\frac{1}{q_{i}}}\lVert u_{j}\rVert_{q_{0}}$
whose right-hand side is finite since
$\sum_{i=0}^{\infty}\frac{1}{q_{i}}=\frac{1}{q_{0}}\sum_{i=0}^{\infty}\frac{1}{\delta_{i}}<\infty\quad\text{and}\quad\sum_{i=0}^{\infty}\frac{\log q_{i}}{q_{i}}<\infty.$
Therefore, letting $k\to\infty$, we can conclude.
Step 3. The continuity of the solutions $u_{j}$ comes from the non-homogeneous Harnack inequality for $X$-elliptic operators established in [22, Theorem 5.5] combined with some standard arguments, see for example [23, Corollary 4.18] for further details. ∎
In the next two lemmas, we prove that the solution $\{u_{a}\}$ is uniformly bounded from below over several spaces.
Lemma 5.2.
Let $a_{1}$ as in Theorem 4.3. Then there exists $C_{1}>0$ such that $\lVert u_{a}\rVert_{\gamma}\geq C_{1}$ for all $a\in(0,a_{1})$.
Proof.
We notice that $F_{a}(t)\geq-a\lvert t\rvert$ for all $t\in\mathbb{R}$. For $\delta$ as given in Theorem 4.3 and from (17) we write $I_{a}(u_{a})\geq\delta$, for all $a\in(0,a_{1})$. Using the embedding of $D_{\gamma}(\mathbb{R}^{N})$ proved in Lemma 3.1 we have
$\delta\leq I_{a}(u_{a})\leq\frac{1}{2}\lVert u_{a}\rVert_{\gamma}^{2}+a\int_{\mathbb{R}^{N}}g(z)\lvert u_{a}(z)\rvert\,{\mathrm{d}z}\leq\frac{1}{2}\lVert u_{a}\rVert_{\gamma}^{2}+a_{1}C\lVert u_{a}\rVert_{\gamma}.$
So there exists $C_{1}=C_{1}(N,\gamma,g,a_{1},\delta)$ such that $\lVert u_{a}\rVert_{\gamma}\leq C_{1}$ for all $a\in(0,a_{1})$. ∎
In what follows, we show an estimate from below to the norm $L^{\infty}(\mathbb{R}^{N})$ of $u_{a}$ for $a$ small enough.
Lemma 5.3.
There exists $a_{3}\in(0,a_{2})$ and $\beta>0$ that does not depend on $a\in(0,a_{3})$, such that $\lVert u_{a}\rVert_{\infty}\geq\beta$ for all $a\in(0,a_{3})$.
Proof.
For the constant $\delta$ given in Theorem 4.3, from (17) we write $I_{a}(u_{a})\geq\delta$ for all $a\in(0,a_{1})$. Further, we have (notice again that $F_{a}(t)\geq-a\lvert t\rvert$ for all $t\in\mathbb{R}$),
$\frac{\lVert u_{a}\rVert_{\gamma}^{2}}{2}=I_{a}(u_{a})+\int_{\mathbb{R}^{N}}g(z)F_{a}(u_{a})\,{\mathrm{d}z}\geq\delta-a\int_{\mathbb{R}^{N}}g(z)\lvert u_{a}\rvert\,{\mathrm{d}z}.$
For every $a\in(0,a_{1})$, using the continuous embedding $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{1}(\mathbb{R}^{N},\lvert g\rvert)$, Lemma 3.1, with embedding constant $C_{2}=C_{2}(N,\gamma,g)$ and the uniform boundeness of $\{u_{a}\}$ in $D_{\gamma}(\mathbb{R}^{N})$, Theorem 4.3, we obtain
$\frac{\lVert u_{a}\rVert_{\gamma}^{2}}{2}\geq\delta-aC_{2}\lVert u_{a}\rVert_{\gamma}\geq\delta-aC_{2}C=\delta-aC_{3},$
where $C_{3}=C_{2}C$. Now if we choose $a_{3}$ such that $0<a_{3}<\min\left\{\frac{\delta}{C_{3}},a_{2}\right\}$, then
$\frac{\lVert u_{a}\rVert_{\gamma}^{2}}{2}\geq\delta_{0}=\delta-a_{3}C_{3}>0,\quad\forall a\in(0,a_{3}).$ (28)
Since $u_{a}$ is a weak solution of (SP), then
$\int_{\mathbb{R}^{N}}\nabla_{\gamma}u_{a}\nabla_{\gamma}\varphi\,{\mathrm{d}z}=\int_{\mathbb{R}^{N}}g(z)f_{a}(u_{a})\varphi\,{\mathrm{d}z},\quad\forall\varphi\in D_{\gamma}(\mathbb{R}^{N}).$
For $\varphi=u_{a}$ we have
$\int_{\mathbb{R}^{N}}\lvert\nabla_{\gamma}u_{a}\rvert^{2}\,{\mathrm{d}z}=\int_{\mathbb{R}^{N}}g(z)f_{a}(u_{a})u_{a}\,{\mathrm{d}z}$
So, by the estimation found in (28)
$\int_{\mathbb{R}^{N}}g(z)f_{a}(u_{a})u_{a}\,{\mathrm{d}z}\geq 2\delta_{0}>0,\quad\forall a\in(0,a_{3}).$
Thus, (2) gives
$2\delta_{0}\leq\int_{\mathbb{R}^{N}}g(z)\left(C\left(\lvert u_{a}\rvert^{\alpha-1}+\lvert u_{a}\rvert\right)+a\right)\,{\mathrm{d}z}\leq\left(C\left(\lVert u_{a}\rVert_{\infty}^{\alpha-1}+\lVert u_{a}\rVert_{\infty}\right)+a\right)\lVert g\rVert_{1}.$
This implies that $\lVert u_{a}\rVert_{\infty}\geq\beta$ for some $\beta>0$ for all $a\in(0,a_{3})$, taking $a_{3}$ smaller if necessary. ∎
6 Positive solutions
As in the previous Section, given a sequence $a_{j}\to 0$ as $j\to\infty$, let us consider the solution $u_{j}=u_{a_{j}}$ for (SP). We denote by $u$ the limit of $u_{j}$ as in Lemma 5.1.
Now consider the function $f_{0}$
$f_{0}(t)=\begin{cases}f(t),\quad&\hbox{if $t\geq 0$};\\ 0,\quad&\hbox{if $t<0$}.\end{cases}$
Proposition 6.1.
$u$ is a weak solution of
$-\Delta_{\gamma}u=g(z)f_{0}(u)\quad\hbox{in $\mathbb{R}^{N}$}.$ (29)
Moreover, $u\in L^{p}(\mathbb{R}^{N})\cap C(\mathbb{R}^{N})$ for every $p\in[2^{*}_{\gamma},\infty]$.
Proof.
Since $a_{j}\to 0$, there exists $j_{1}\in\mathbb{N}$ such that $a_{j}\leq a_{2}$ for all $j\geq j_{1}$ (with $a_{2}$ we denote the parameter $a_{2}$ as in the statement of Lemma 5.1). Now, for every $\varphi\in D_{\gamma}(\mathbb{R}^{N})$ with $\varphi\geq 0$ we show that
$\int_{\mathbb{R}^{N}}g(z)f_{j}(u_{j})\varphi(z)\,{\mathrm{d}z}\to\int_{\mathbb{R}^{N}}g(z)f_{0}(u)\varphi(z)\,{\mathrm{d}z},\quad\hbox{as $j\to\infty$}.$
We split
$\lvert f_{j}(u_{j})-f_{0}(u)\rvert\leq\lvert f_{j}(u_{j})-f_{0}(u_{j})\rvert+\lvert f_{0}(u_{j})-f_{0}(u)\rvert.$
By the continuity of $K_{0}^{\prime}$,
$\int_{\mathbb{R}^{N}}g(z)\lvert f_{0}(u_{j})-f_{0}(u)\rvert\varphi(z)\,{\mathrm{d}z}\to 0,\quad\hbox{as $j\to\infty$}.$
Further, $\lvert f_{j}(u_{j})-f_{0}(u_{j})\rvert\leq a_{j}$. Therefore,
$\int_{\mathbb{R}^{N}}g(z)\lvert f_{j}(a_{j})-f_{0}(u)\rvert\varphi(z)\,{\mathrm{d}z}\\ \leq a_{j}\int_{\mathbb{R}^{N}}g(z)\varphi(z)\,{\mathrm{d}z}+\int_{\mathbb{R}^{N}}g(z)\lvert f_{0}(u_{j})-f_{0}(u)\rvert\varphi(z)\,{\mathrm{d}z}\to 0\quad\hbox{as}\leavevmode\nobreak\ j\to\infty.$
From the weak formulation
$\int_{\mathbb{R}^{N}}\nabla_{\gamma}u_{j}(z)\nabla_{\gamma}\varphi(z)\,{\mathrm{d}z}=\int_{\mathbb{R}^{N}}g(z)f_{j}(u_{j})\varphi(z)\,{\mathrm{d}z},\forall\varphi\in D_{\gamma}(\mathbb{R}^{N}),\quad\varphi\geq 0.$
Taking the limit as $j\to\infty$ gives
$\int_{\mathbb{R}^{N}}\nabla_{\gamma}u(z)\nabla_{\gamma}\varphi(z)\,{\mathrm{d}z}=\int_{\mathbb{R}^{N}}g(z)f(u)\varphi(z)\,{\mathrm{d}z},\forall\varphi\in D_{\gamma}(\mathbb{R}^{N}),\quad\varphi\geq 0.$
Now, for any $\varphi\in D_{\gamma}(\mathbb{R}^{N})$, we write $\varphi=\varphi^{+}-\varphi^{-}$. We see that the above identity holds for both $\varphi^{+}$ and $\varphi^{-}$, so
$\int_{\mathbb{R}^{N}}\nabla_{\gamma}u(z)\nabla_{\gamma}\varphi(z)\,{\mathrm{d}z}=\int_{\mathbb{R}^{N}}g(z)f(u)\varphi(z)\,{\mathrm{d}z},\quad\forall\varphi\in D_{\gamma}(\mathbb{R}^{N}),$ (30)
that is, $u$ is a weak solution of (29). With the same argument of Lemma 5.1, we obtain $u\in L^{p}(\mathbb{R}^{N})$, $p\in[2^{*}_{\gamma},\infty]$ and $u$ is a continuous function in $\mathbb{R}^{N}$. ∎
Proposition 6.2.
The function $u$ is positive.
Proof.
Taking the test function $\varphi=u^{-}=\min\left\{u,0\right\}$ in (30) we find that $\lVert u^{-}\rVert_{\gamma}=0$, so $u\geq 0$ a.e. in $\mathbb{R}^{N}$.
Let $\Sigma=\{0\}\times\mathbb{R}^{\ell}$ be the degenerate set for the Grushin operator. If $\Omega$ is any bounded domain contained in $\mathbb{R}^{N}\setminus\Sigma$, then $\Delta_{\gamma}$ is a uniformly elliptic operator on $\Omega$. Hence $u\in C^{2}(\Omega)$ and either $u\equiv 0$ on $\Omega$ or $u>0$ in $\Omega$. In the former case, it follows from [26, Proposition 2.5] that $u\equiv 0$ on the connected component of $\mathbb{R}^{N}\setminus\Sigma$ which contains $\Omega$. Therefore $u\equiv 0$ on $\mathbb{R}^{N}\setminus\Sigma$.
Now, it is clear that $\Sigma$ has empty interior, and it follows from [25, Theorem 9, Chapter 1] that $\mathbb{R}^{N}\setminus\Sigma$ is dense in $\mathbb{R}^{N}$. Recalling that $u$ is a continuous function, we conclude that $u\equiv 0$ on $\Sigma$. To this stage the function $u$ might be zero in (the interior of) some connected component of $\mathbb{R}^{N}\setminus\Sigma$ and strictly positive in (the interior of) some other component. However this is again impossible, since [26, Proposition 2.5] applies again on any ball which meets two such components across $\Sigma$. Hence either $u\equiv 0$ on $\mathbb{R}^{N}$, or $u>0$ on $\mathbb{R}^{N}$. The former is not admissible due to the Lemma 5.3. ∎
Remark 6.3.
When $\gamma$ is an even integer, a very elegant proof of the positivity of $u$ can be devised. We follow an idea of [6]. When $\gamma$ is an integer, it is known that $-\Delta_{\gamma}$ satisfies the Hörmander condition. Recalling that $g>0$ and $f_{0}\geq 0$, we see that $\int_{\mathbb{R}^{N}}\nabla_{\gamma}u\nabla_{\gamma}\varphi\,{\mathrm{d}z}\geq 0$ for every $\varphi\in C_{0}^{\infty}(\mathbb{R}^{N})$ such that $\varphi\geq 0$. Hence $\int_{\mathbb{R}^{N}}u(-\Delta_{\gamma}\varphi)\,{\mathrm{d}z}\geq 0$ for every such $\varphi$. Recalling that $u$ is a continuous function, it follows that $u$ is a viscosity subsolution of the operator $-\Delta_{\gamma}$. The strong maximum principle proved in [4, Corollary 1.4] yields that either $u\equiv 0$ or $u>0$ in $\mathbb{R}^{N}$. The first option, again, is not admissible due to the Lemma 5.3.
Now we are going to show that $u_{j}\to u$ in $L^{\infty}(\{0\}\times\mathbb{R}^{\ell})$. We essentially follow [1, 7], although a few important differences will arise.
For $z=(x,y)\in\mathbb{R}^{N}=\mathbb{R}^{m}\times\mathbb{R}^{\ell}$, we denote by
$d(z)=\left(\lvert x\rvert^{2\left(\gamma+1\right)}+(\gamma+1)^{2}\lvert y\rvert^{2}\right)^{\frac{1}{2\left(\gamma+1\right)}}$ (31)
the homogeneous norm associated to $\Delta_{\gamma}$. Note that this norm is homogeneous of degree $1$ with respect to the anisotropic dilatation
$\delta_{\lambda}(x,y)=(\lambda x,\lambda^{\gamma+1}y),\quad\lambda>0.$
and, for $\gamma=0$, it reduces to the usual Euclidean distance in $\mathbb{R}^{N}$.
The function
$\Gamma(z)=\frac{C}{d(z)^{N_{\gamma}-2}}$
is a fundamental solution for the Grushin operator with singularity at $z=0$, see [5] or [16, Appendix B].
The constant $C$ is a suitable positive constant that depends on $m,\ell$ and $\gamma$. Precisely,
$C^{-1}=(N_{\gamma}-2)\int\limits_{\{z\mid d(z)=1\}}\frac{\lvert x\rvert^{2\gamma}}{\left(\lvert x\rvert^{2\left(1+2\gamma\right)}+\left(1+\gamma\right)^{2}\lvert y\rvert^{2}\right)^{1/2}}\,\mathrm{d}S,$
where $S$ denotes the surface measure of the manifold $\left\{z\mid d(z)=1\right\}$.
We recall the following definition.
Definition 6.4.
A function $u\colon\mathbb{R}^{N}\to\mathbb{R}$ is radially symmetric (radial for short) if there exists a function $v\colon\mathbb{R}\to\mathbb{R}$ such that
$u(z)=v(r)$
where $r=d(z)$, for all $z\in\mathbb{R}^{N}$.
For $R>0$ we set $B_{R}=\left\{\xi\in\mathbb{R}^{N}\mid d(\xi)<R\right\}$.
Let us consider a spherical shell $\Omega=B_{R_{2}}\setminus\overline{B_{R_{1}}}$, with $0\leq R_{1}<R_{2}\leq+\infty$. If $u$ is a radial function, by [15, Eq. (2.4)] we get the formula
$\int_{\Omega}u(d(\xi))\,\mathrm{d}\xi=s_{N}\int_{R_{1}}^{R_{2}}\rho^{N_{\gamma}-1}u(\rho)\,\mathrm{d}\rho,$ (32)
where $s_{N}$ is an explicit constant that depends only on $N$.
Proposition 6.5.
$\lVert u_{j}-u\rVert_{L^{\infty}(\{0\}\times\mathbb{R}^{\ell})}\to 0$ as $j\to\infty$.
Proof.
By [18, Sect. 3.1] and recalling that $u$ is a continuous function, we have the representation formula
$u(z)=C\int_{\mathbb{R}^{N}}\frac{g(\xi)f_{0}(u(\xi))}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\quad\hbox{for a.e. $z\in\{0\}\times\mathbb{R}^{\ell}$}.$ (33)
Moreover, since $u_{j}$ is a solution of (SP), we also have
$u_{j}(z)=C\int_{\mathbb{R}^{N}}\frac{g(\xi)f_{j}(u_{j}(\xi))}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\quad\hbox{for a.e. $z\in\{0\}\times\mathbb{R}^{\ell}$}.$ (34)
Using (33) and (34) we estimate $\lvert u-u_{j}\rvert$ as follows: for every $z\in\{0\}\times\mathbb{R}^{\ell}$,
$\lvert u(z)-u_{j}(z)\rvert\leq C\int_{\mathbb{R}^{N}}g(\xi)\frac{\lvert f_{j}(u_{j}(\xi))-f_{0}(u(\xi))\rvert}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\\ \leq C\left(\int_{B_{1}}g(\xi)\frac{\lvert f_{j}(u_{j}(\xi))-f_{0}(u(\xi))\rvert}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi+\int_{\mathbb{R}^{N}\setminus B_{1}}g(\xi)\frac{\lvert f_{j}(u_{j}(\xi))-f_{0}(u(\xi))\rvert}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\right).$ (35)
Take $\delta>1$. Applying the Hölder’s inequality with the conjugate pairs $(\delta,\tilde{\delta})$ we estimate the first integral of (35) as
$\int_{B_{1}}g(\xi)\frac{\left|f_{j}(u_{j}(\xi))-f_{0}(u(\xi))\right|}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\leq\left(\int_{B_{1}}\frac{1}{d(z-\xi)^{(N_{\gamma}-2)\delta}}\,\mathrm{d}\xi\right)^{\frac{1}{\delta}}\times\\ \left(\int_{B_{1}}g(\xi)^{\tilde{\delta}}\left|f_{j}(u_{j}(\xi))-f_{0}(u(\xi))\right|^{\tilde{\delta}}\,\mathrm{d}\xi\right)^{\frac{1}{\tilde{\delta}}}.$ (36)
We compute, using (32)
[…]
Moreover, the second term in the inequality (36) converges to zero. In fact, proceeding as in Lemma 5.1, $u_{j}\to u$ in $L^{p},\,p\in[2^{*}_{\gamma},\infty)$. Then we can apply the Dominated Convergence Theorem.
Next, the second integral of (36) has the following bound:
$\int_{\mathbb{R}^{N}\setminus B_{1}}g(\xi)\frac{\lvert f_{j}(u_{j}(\xi))-f_{0}(u(\xi))\rvert}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\leq\int_{\mathbb{R}^{N}\setminus B_{1}}g(\xi)\lvert f_{j}(u_{j}(\xi))-f_{0}(u(\xi))\rvert\,\mathrm{d}\xi.$
Again, by the generalized dominated convergence theorem,
$\lim_{j\to+\infty}\int_{\mathbb{R}^{N}\setminus B_{1}}g(\xi)\frac{\lvert f_{j}(u_{j}(\xi))-f_{0}(u(\xi))\rvert}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi=0.$
Therefore, (35) yields $u_{j}\to u$ in $L^{\infty}(\{0\}\times\mathbb{R}^{\ell})$ as $j\to\infty$. ∎
The immediate corollary concludes the proof of the first statement of Theorem 1.1.
Corollary 6.6.
$u_{j}\geq 0$ a.e. in $\{0\}\times\mathbb{R}^{\ell}$ for $j$ large enough.
Let us now conclude the proof of Theorem 1.1.
Theorem 6.7.
If (f3) and (g1) are satisfied then $u_{j}>0$ a.e. in $\{0\}\times\mathbb{R}^{\ell}$.
Proof.
For each $j\in\mathbb{N}$, since $f_{j}$ is locally Lipschitz (f3) and $0\leq u_{j},u\leq C$, we have
$\lvert f_{j}(u_{j}(z))-f_{0}(u(z))\rvert\leq M\lvert u_{j}(z)-u(z)\rvert,$
for some $M>0$. For $z\in\{0\}\times\mathbb{R}^{\ell}\setminus\{(0,0)\}$, using (33) and (34), we get
$\lvert u_{j}(z)-u(z)\rvert\leq C\left(M\int_{\mathbb{R}^{N}}g(\xi)\frac{\lvert u_{j}(\xi)-u(\xi)\rvert}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi+a_{j}\int_{\mathbb{R}^{N}}\frac{g(\xi)}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\right).$
Since $g$ satisfies (g1), from the above inequality and the previous theorem we get
$\lvert u_{j}(z)-u(z)\rvert\leq C\left(M\lVert u_{j}-u\rVert_{\infty}+a_{n}\right)\frac{C(g)}{d(z)^{N_{\gamma}-2}}.$
Hence
$\sup_{z\in\mathbb{R}^{N}\setminus\{0\}}d(z)^{N_{\gamma}-2}\lvert u_{j}(z)-u(z)\rvert\to 0\leavevmode\nobreak\ \hbox{as $j\to\infty$}.$ (37)
We claim that
$\lim_{d(z)\to+\infty}d(z)^{N_{\gamma}-2}u(z)>0.$ (38)
Using (33) it follows that
$\lim_{d(z)\to+\infty}d(z)^{N_{\gamma}-2}u(z)=C\lim_{d(z)\to+\infty}\int_{\mathbb{R}^{N}}\frac{g(\xi)f_{0}(u(\xi))d(z)^{N_{\gamma}-2}}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\\ \geq C\int_{B_{R}}\frac{g(\xi)f_{0}(u(\xi))d(z)^{N_{\gamma}-2}}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi$ (39)
for any $R>0$. For any $R>0$ there exists $z\in\{0\}\times\mathbb{R}^{\ell}$ such that $d(z)>2R+1$. Hence
$d(z-\xi)^{N_{\gamma}-2}\geq\lvert d(z)-d(\xi)\rvert^{N_{\gamma}-2}\geq\lvert d(z)-R\rvert^{N_{\gamma}-2}\geq 2^{2-N_{\gamma}}\left(1+d(z)\right)^{N_{\gamma}-2},$
for every $\xi\in B_{R}$. Using the above estimate, for $\xi\in B_{R}$ we get
$\frac{g(\xi)f_{0}(u(\xi))d(z)^{N_{\gamma}-2}}{d(z-\xi)^{N_{\gamma}-2}}\leq 2^{N_{\gamma}-2}g(\xi)f_{0}(u(\xi)).$
Furthermore,
$\frac{g(\xi)f_{0}(u(\xi))d(z)^{N_{\gamma}-2}}{d(z-\xi)^{N_{\gamma}-2}}\to g(\xi)f_{0}(u(\xi))$
a.e. in $B_{R}$, as $d(z)\to\infty$. Therefore, the Dominated Convergence Theorem yields
$\lim_{d(z)\to+\infty}\int_{B_{R}}\frac{g(\xi)f(u(\xi))d(z)^{N_{\gamma}-2}}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi=\int_{B_{R}}g(\xi)f_{0}(u(\xi))\,\mathrm{d}\xi.$
So, from (39) we conclude that
$\lim_{d(z)\to+\infty}d(z)^{N_{\gamma}-2}u(z)\geq C\int_{B_{R}}g(\xi)f_{0}(u(\xi))\,\mathrm{d}\xi.$
Letting $R\to\infty$ and applying the Fatou’s lemma we get
$\lim_{d(z)\to\infty}d(z)^{N_{\gamma}-2}u(z)\geq C\int_{\mathbb{R}^{N}}g(\xi)f_{0}(u(\xi))\,\mathrm{d}\xi.$
So (38) follows by (33) and Proposition 6.2.
Therefore, by (37) there exists $j_{2}\in\mathbb{N}$ and $R>>1$ such that for $j\geq j_{2},\,u_{j}>0$ a.e. on ($(\{0\}\times\mathbb{R}^{\ell})\setminus B_{R}$. Moreover, since $u$ is continuous, there exists $\eta>0$ such that $u>\eta$ on $\overline{B_{R}}$. Therefore, from Proposition 6.5, there exists $j_{3}\in\mathbb{N}$ such that for $j\geq j_{3},\,u_{j}>0$ a.e. on $\overline{B_{R}}\cap(\{0\}\times\mathbb{R}^{\ell})$. Thus, by choosing $j_{4}=\max\left\{j_{2},j_{3}\right\}$, we see that for $j\geq j_{4},\,u_{j}>0$ a.e. on $\{0\}\times\mathbb{R}^{\ell}$. This completes the proof. ∎
Remark 6.8.
The validity of the previous argument relies heavily on the representation formula (33), which is typically derived from the existence of a Green function associated to the differential operator. For linear differential operators with constant coefficients, the existence of a fundamental solution allows one to construct (continuous) solutions via convolution. by exploiting the invariance of the operator under translations. The Grushin operator $\Delta_{\gamma}$ is a differential operator with variable coefficients, and in particular it is invariant under translations only in the subspace $\{0\}\times\mathbb{R}^{\ell}$ of $\mathbb{R}^{N}$. So, even if a fundamental solution with singularity at a fixed point is available, it is not trivial to use it in order to derive an explicit Green function. As we have seen, it was proved in [18] that (33) holds true at points of the form $(0,y)$ at which the weak solution $u$ is continuous. It is interesting to observe that the authors of [17] proved the equivalence of a representation formula like (33) and of a Liouville-type result for the operator $\Delta_{\gamma}$, but only for $C^{2}$ solutions. However a full regularity theory for $\Delta_{\gamma}$ is not available, this approach remains unapplicable.
Furthermore, the existence of an approximate Green function for $\Delta_{\gamma}$ was proved in [29] and exploited in [2]. Unfortunately the known properties of it are too weak for our purposes, and would lead to replace (g1) with a completely implicit assumption.